Che Kubin Companys relevant range of production is 29,000 to 33,000 units. When it produces and sells 31.000 units, its average costs per unit are as follows: Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Fixed selling expense Fixed administrative expense Sales commissions Variable administrative expense Average Cost per Unit 8.90 $5.90 $ 3.40 $ 6.90 $ 5.40 4.40 $ 2.90 $2.40 Required 1. For financial accounting purposes, what is the total amount of product costs incurred to make 31,000 units? 2. For financial accounting purposes, what is the total amount of period costs incurred to sell 31,000 units? 3. For financial accounting purposes, what is the total amount of product costs incurred to make 33,000 units? 4. For financial accounting purposes, what is the total amount of period costs incurred to sell 29,000 units? (For all requirements, do not round intermediate calculations.)
Required: 1. For financial accounting purposes, what is the total amount of product costs incurred to make 31,000 units? 2. For financial accounting purposes, what is the total amount of period costs incurred to sell 31,000 units? 3. For financial accounting purposes, what is the total amount of product costs incurred to make 33,000 units? 4. For financial accounting purposes, what is the total amount of period costs incurred to sell 29,000 units? For all requirements, do not round intermediate calculations.) 1. Total amount of product costs 2. Total amount of period costs incurred 3. Total amount of product costs 4. Total amount of period costs
Answer #1
1
Direct materials 275900 =31000*8.9
Direct labor 182900 =31000*5.9
Variable manufacturing overhead 105400 =31000*3.4
Fixed manufacturing overhead 213900 =31000*6.9
Total amount of product costs 778100
2
Fixed selling expense 167400 =31000*5.4
Fixed administrative expense 136400 =31000*4.4
Sales Commissions 89900 =31000*2.9
Variable administrative expense 74400 =31000*2.4
Total amount of period costs incurred 468100
3
Direct materials 293700 =33000*8.9
Direct labor 194700 =33000*5.9
Variable manufacturing overhead 112200 =33000*3.4
Fixed manufacturing overhead 213900 =31000*6.9
Total amount of product costs 814500
4
Fixed selling expense 167400 =31000*5.4
Fixed administrative expense 136400 =31000*4.4
Sales Commissions 84100 =29000*2.9
Variable administrative expense 69600 =29000*2.4
Total amount of period costs 457500